AO Retroreflector Work: 21 October 2010 UT

Observer: DMiller, SEsposito, MXompero, LBusoni, APuglisi, FQuiros
Telescope Operator: DHuerta, GBechetti (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: TShih (LBTO)

Summary

Rain and clouds kept us closed all night, so retroreflector work proceeded through the night

We were able be keep he loop closed while moving the telescope up and down in elevation. We move in small increments (1 and 3 degree steps) from 90 to 26 degrees and then back to 90. Mirror positions were recorded and post-processed to determine the amount of asigmatism induced as a function of elevation. Two beautiful curves were measured, up and down. These curves were not the same, so there is obvious and large amounts of asigmastism hysteresis. Marco is working on the data.

On Thursday morning Marco told me the Astigmatism at 29 elevation was ~7 micron rms. If I recall this is lower than the maximum we saw in the Spring of 10-12 microns.

Preparations

Details

-- DougMiller - 21 Oct 2010
Topic revision: r3 - 27 Oct 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback